OpenStudy (anonymous):

What type of conic section is the following equation? x2 - 4y2 = 100

OpenStudy (solomonzelman):

\(\normalsize\color{black}{ x^2-4y^2=100 }\) \(\normalsize\color{black}{ \frac{\LARGE{x^2} }{\LARGE{100} }-\frac{\LARGE{y^2} }{\LARGE{25} }=1 }\)

OpenStudy (anonymous):

@SolomonZelman Ellipse, circle, parabola?